# Adaptive Deep Learning for Multiclass Breast Cancer Classification via Misprediction Risk Analysis

### v1

@@ 6.3 Evaluation Results of Risk-based Adaptive Learning
We have also visualized the comparative results by confusion matrix as shown in Figure 7. In Figure 7 (a), it can be observed that MultiRisk correctly classifies 364 of 570 test images on the original BRACS dataset, achieving the highest accuracy in ”Invasive Carcinoma” (75/81), while mispredictions mainly occurred between ”Pathological Benign,” ”Usual Ductal Hyperplasia,” and ”Atypical Ductal Hyperplasia.” In Figure 7 (b), for the BRACS(512 x 512) dataset, MultiRisk accurately predicts 375 of 512 images, with similar errors in distinguishing between ”Pathological Benign,” ”Usual Ductal Hyperplasia,” and ”Atypical Ductal Hyperplasia.” Finally, for the domain transfer analysis, the confusion matrix in Figure 7 (c) shows that, using 80 test images from BRACS-BACH, MultiRisk accurately classifies 64 images, with mispredictions mainly between ”Normal” and ”Benign” classes, highlighting both strengths and areas for improvement in class differentiation and domain transfer.
@@
6.4 Sensitivity Study of Risk Analysis
In the sensitivity study, we have evaluated the sensitivity of risk model w.r.t validation samples by conducting two experiments, the first experiment using a different number of validation samples, and the second one using the same number but randomly selected validation samples. Note that the training and test samples remain the same in the sensitivity study. Additionally, we have also evaluated the efficacy of MultiRisk by leveraging different baseline classifier models.
@@ 6.4 Sensitivity Study of Risk Analysis
Furthermore, the evaluation results on the performance sensitivity of MultiRisk-based adaptive learning with different Baseline models have been presented in Table 11. It can be observed that on each baseline model, MultiRisk outperforms both the baseline and LearnRisk. These results clearly demonstrate the performance robustness of the MultiRisk-based approach.
@@
Accuracy F1 Accuracy F1 Accuracy F1
Resnet50 63.33 62.88 64.21 63.40 65.04 64.08
EfficientNetB4 58.25 56.74 58.95 56.36 59.47 57.03
CCT 60.7 60.27 61.58 60.27 62.90 61.97
Densenet201 63.16 62.97 63.33 63.03 63.41 63.1
Densenet121 63.33 63.44 65.09 64.23 66.34 65.98
@@ 6.5 Ablation Study
To validate the design choices made in MultiRisk and MultiRisk-based adaptive learning, we have conducted an ablation study, in which we remove certain new mechanisms during each major step and compare the performance of combined MultiRisk Mechanism with selected original LearnRisk based mechanisms.
Table 12: Comparison of Methods on Different Datasets where each time we run the approach excluding one step from our method and including the method from LearnRisk
Method Dataset Accuracy F1 Score Risk AUROC BRACS(Original) 59.65 ± 2.0 58.44 ± 1.8 73.2 Risk Feature Generation (LearnRisk) BRACS(512 x 512) 63.86 ± 3.8 63.41 ± 2.4 75.3 BRACS-BACH 75.00 ± 12.1 75.63 ± 10.3 67.8 Risk Model Construction/Training (LearnRisk) BRACS(Original) 62.81 ± 1.2 61.20 ± 1.7 78.1 BRACS(512 x 512) 65.07 ± 3.7 64.23 ± 4.9 76.2 BRACS-BACH 75.00 ± 9.3 75.63 ± 10.3 72.0 Risk Based Adaptive Training (LearnRisk) BRACS(Original) 61.40 ± 4.3 59.43 ± 5.1 73.5 BRACS(512 x 512) 63.86 ± 1.3 62.87 ± 2.7 75.3 BRACS-BACH 80.00 ± 1.4 80.53 ± 1.1 69.8 MultiRisk (Ours) BRACS(Original) 63.51 ± 2.0 61.15 ± 1.4 78.1 BRACS(512 x 512) 66.34 ± 2.3 65.98 ± 1.0 75.6 BRACS-BACH 80.00 ± 1.3 80.53 ± 0.9 76.3
The detailed evaluation results in terms of ROC have been presented in Table 12. It can be observed that MultiRisk achieves the overall highest performance at 78.1%, and 76.3% on the original BRACS, and BRACS-BACH datasets respectively, while a nominal difference was recorded for BRACS(512x512) dataset where MultiRisk still performs the second best. The evaluation results in terms of accuracy and F1 have also been shown in Table 12. In most cases, MultiRisk outperforms other methods, although for the domain transfer task, accuracy and F1 scores are often similar due to the fewer validation and test images compared to training images.
@@ 7 Conclusion and Future Work
In this paper, we have proposed a novel approach to misprediction risk analysis, MultiRisk, for multiclass breast cancer detection. MultiRisk introduces new mechanisms for risk feature generation, risk model construction and training, and risk-based adaptive training, which are specifically designed for multiclass problems and effectively enable accurate misprediction quantification. Based on the MultiRisk approach, we have also presented an adaptive learning solution that can efficiently fine-tune a baseline multiclass model to a target workload. Our extensive experiments on real benchmark datasets validate the efficacy of the proposed solutions.
For future work, it is noteworthy that the proposed MultiRisk approach is potentially applicable to other types of cancer detection; however, its technical details require further investigation. Additionally, the current solution extracts risk metrics based on mainstream deep neural network (DNN) models, which have limited overall interpretability. To enhance both the efficacy and interpretability of risk analysis, future research could explore generating risk features using explainable AI techniques.
@@
Gul Sheeraz led the study, handled data curation, formal analysis, investigation, methodology development, project administration, and drafting. Qun Chen provided funding, resources, supervision, and manuscript review. Liu Feiyu contributed to analysis, investigation, methodology, software, validation, visualization, and manuscript review. Zhou Fengjin provided supervision related to medical information for this task.

### v2

@@ 6.3 Evaluation Results of Risk-based Adaptive Learning
We have also visualized the comparative results by confusion matrix as shown in Figure 6. In Figure 6 (a), it can be observed that MultiRisk correctly predicts 364 of 570 test images on the original BRACS dataset, achieving the highest accuracy in "Invasive Carcinoma" (75/81), while mispredictions mainly occurred between "Pathological Benign," "Usual Ductal Hyperplasia," and "Atypical Ductal Hyperplasia." In Figure 6 (b), for the BRACS(512 x 512) dataset, MultiRisk accurately predicts 375 of 512 images, with similar errors in distinguishing between "Pathological Benign," "Usual Ductal Hyperplasia," and "Atypical Ductal Hyperplasia." Finally, for the domain transfer analysis, the confusion matrix in Figure 6 (c) shows that, using 80 test images from BRACS-BACH, MultiRisk accurately classifies 64 images, with mispredictions mainly between "Normal" and "Benign" classes, highlighting both strengths and areas for improvement in class differentiation and domain transfer.
@@
6.4 Sensitivity Analysis
In the sensitivity analysis, we have evaluated the sensitivity of risk model w.r.t validation samples by conducting two experiments, the first experiment using a different number of validation samples, and the second one using the same number but randomly selected validation samples. Note that the training and test samples remain the same in the sensitivity study. Additionally, we have also evaluated the efficacy of MultiRisk by leveraging different baseline prediction models.
@@ 6.4 Sensitivity Analysis
Furthermore, the evaluation results on the performance sensitivity of MultiRisk-based adaptive learning with different Baseline models have been presented in Table 9. It can be observed that on each baseline model, MultiRisk outperforms both the baseline and LearnRisk in terms of f1 and AUC. These results clearly demonstrate the performance robustness of the MultiRisk-based approach.
@@
F1 AUC F1 AUC F1 AUC
Resnet50 62.88 89.64 63.40 88.93 64.08 91.04
EfficientNetB4 56.74 89.17 56.36 89.95 57.03 91.49
CCT 60.27 89.13 60.27 90.61 61.97 92.03
Densenet201 62.97 89.98 63.03 90.31 63.10 91.48
Densenet121 63.44 90.73 64.23 90.10 65.98 92.07
@@ 6.5 Ablation Study
To validate the design choices in Multiclass misprediction risk analysis and the MultiRisk-based adaptive learning framework, we conducted an ablation study comparing four key components. First, for risk feature generation, we evaluated the impact of using features from each model without feature selection and fusion. Second, to assess the Risk Model, we compared our attention-based weighting mechanism with traditional Gaussian weighting. Third, we evaluated our voting-based learning-to-rank strategy against conventional pairwise comparison-based ranking. Finally, to examine the handling of class imbalance, we compared our approach with established methods such as Focal Loss and Class-Balanced Loss. For feature extraction. attention based weighting, and learning-to-rank experiments, F1 scores were recorded for adaptive learning and Risk-AUROC for risk analysis. For class imbalance assessment, both F1 and AUC metrics were computed. Across all scenarios, MultiRisk consistently outperformed alternative configurations. The results are summarized in Table 10.
Table 10: Ablation Study: Comparative Analysis of Our Design Choices and Innovations with Traditional Methods
Feature Extraction
Method BRACS BRACS-512 BRACS-BACH
F1 (%) Risk-AUROC (%) F1 (%) Risk-AUROC (%) F1 (%) Risk-AUROC (%)
Resnet50 59.03 70.1 62.35 69.2 76.12 65.8
Densenet121 60.86 72.2 64.81 73.9 76.56 67.4
EfficientnetB4 59.87 65.7 62.29 70.2 77.86 73.3
CCT 53.93 70.4 62.18 74.7 72.81 72.3
MultiRisk(Ours) 61.15 78.1 65.98 75.6 80.53 75.3
Risk Model Weights
Method BRACS BRACS-512 BRACS-BACH
F1 (%) Risk-AUROC (%) F1 (%) Risk-AUROC (%) F1 (%) Risk-AUROC (%)
Gaussian Weights 60.20 78.1 64.23 75.63 76.12 72.0
MultiRisk(Ours) 61.15 78.1 65.98 75.6 80.53 75.3
Learning to Rank
Method BRACS BRACS-512 BRACS-BACH
F1 (%) Risk-AUROC (%) F1 (%) Risk-AUROC (%) F1 (%) Risk-AUROC (%)
Pairwise Comparison 60.05 76.4 64.23 73.8 77.98 74.3
Voting Based(Ours) 61.15 78.1 65.98 75.6 80.53 75.3
Class Imbalance
Method BRACS BRACS-512 BRACS-BACH
F1 (%) AUC (%) F1 (%) AUC (%) F1 (%) AUC (%)
Focal Loss 60.92 91.82 64.37 91.37 76.75 88.03
Class-Balanced Loss 58.92 89.62 62.35 90.12 78.69 90.85
MultiRisk(Ours) 61.15 91.38 65.98 92.07 80.53 93.23
6.6 Generalization Analysis
We evaluated MultiRisk on two diverse lung-colon cancer LC25000 [98], and lung cancer LungHist700 [98] histopathology datasets to test its generalizability beyond breast cancer. These results are presented in table 11, compared to baseline model(Densenet121), MultiRisk consistently achieves higher F1 and AUC scores across almost all the data scenarios.
Notably, in domain transfer scenarios like training on LC25000 and testing on LungHist700 and vice versa, MultiRisk maintains superior performance, demonstrating strong adaptability. These results highlight MultiRisk’s effectiveness in improving prediction accuracy and robustness across different datasets and cancer types, supporting its broader clinical applicability.
Furthermore, our misprediction risk analysis leverages features extracted from deep learning models, making MultiRisk readily adaptable to various neural network architectures, including complex multi-modal Vision-Language Models (VLMs). To demonstrate our method’s generalization strength, we evaluated it on three supervised architectures: CLIP-style ResNet-50 with GPT-2, ResNet-50 with LLaVA-Med, and ResNet-50 with DeepSeek. As shown in Table 12, MultiRisk delivered substantial performance gains over the baseline VLMs in all scenarios.
Table 11: Generalization Results: Baseline vs. MultiRisk on LC25000 and LungHist700 Datasets
Dataset Baseline MultiRisk
F1 (%) AUC (%) F1 (%) AUC (%)
LC25000 96.34 99.53 98.19 99.24
LungHist700 80.67 92.32 84.88 93.05
LC25000 $\rightarrow$ LungHist700 68.43 90.71 72.64 92.62
LungHist700 $\rightarrow$ LC25000 65.74 90.71 68.73 92.19
Table 12: Generalization Results: Baseline vs. MultiRisk on Vision–Language Models (BRACS Dataset)
Vision–Language Model Baseline MultiRisk
F1 (%) AUC (%) F1 (%) AUC (%)
CLIP(ResNet-50 + GPT2) 51.43 85.21 54.32 86.76
CLIP(ResnNet-50 + LLaVA-Med) 52.11 85.98 54.78 87.85
DeepSeek-VL(ResNet-50) 54.65 87.71 57.41 89.09
6.7 Computational Efficiency and Scalability Discussion
While MultiRisk achieves substantial improvements in prediction performance and robustness, we recognize the critical importance of computational efficiency and scalability in resource-constrained diagnostic settings. To address this, our adaptive learning process is deliberately constrained to a short training duration of only 10 epochs, significantly reducing the computational burden compared to standard prolonged training regimes. Furthermore, we validate MultiRisk on lower-resolution images (512×512), demonstrating competitive and better performance that aligns with the reduced memory and processing requirements inherent in such inputs, making the approach well-suited for deployment on hardware with limited resources.
Additionally, the misprediction risk analysis model is sample-efficient, requiring only a small subset of training samples to effectively calculate risk and guide adaptive learning. This targeted training avoids extensive overhead and contributes to the method’s scalability. Collectively, these design choices ensure that MultiRisk maintains a favorable balance between performance gains and computational practicality, positioning it as a viable solution for real-world clinical applications with constrained computational resources.
@@ 7 Conclusion and Future Work
In this paper, we proposed a novel approach to multiclass misprediction risk analysis, MultiRisk, for breast cancer subtype prediction. MultiRisk introduces new mechanisms for risk feature generation, risk model construction and training, and risk-based adaptive training, all specifically designed for multiclass problems to enable accurate misprediction quantification. Building on MultiRisk, we further presented an adaptive learning strategy capable of efficiently fine-tuning a baseline subtype prediction model to a target workload of breast cancer. Extensive experiments on real benchmark datasets demonstrate the effectiveness, robustness, and generalization ability of our approach across different types of cancer and diverse deep learning models.
For future work, we note that the proposed MultiRisk framework is potentially applicable to other types of prediction tasks in healthcare; however, its technical details warrant further exploration. Additionally, the current solution derives risk metrics from mainstream deep neural network (DNN) models, which inherently have limited interpretability. To improve both the effectiveness and interpretability of risk analysis, future research could investigate generating risk features through explainable AI techniques.
Data Availability
The datasets used in this study are publicly available and open source for research purpose. The BRACS: BReAst Carcinoma Sub-typing can be accessed at https://www.bracs.icar.cnr.it/. ICIAR 2018 Grand Challenge on Breast Cancer Histology Dataset i.e BACH at https://iciar2018-challenge.grand-challenge.org/Dataset/. Additionally, the dataset used for generalization experiments can be accessed at https://zenodo.org/records/14998042 and https://www.kaggle.com/datasets/abdullahhasansajjad/lunghist700. The code for MultiRisk is available at https://github.com/SheerazNWPU/MultiRISK. Additional data or resources can be provided upon reasonable request.
@@
Gul Sheeraz led the study, handled data curation, formal analysis, investigation, methodology development, project administration, and drafting. Qun Chen provided funding, resources, supervision, and manuscript review. Liu Feiyu contributed to analysis, investigation, methodology, software, validation, visualization, and manuscript review. Zhou Fengjin provided co-supervision related to medical information for this task.
@@ Acknowledgments
Conflict of Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.